# A dynamical algorithm to compute hyperbolic Julia sets in polynomial time

### v1

@@ Appendix A More Proof Details
For completeness, we include here more details for this proof. First, applying Corollary 2.19 of [3] to $p^{\nu}$ on $\mathcal{B}^{\prime}_{n}$, we compute a radius $r^{\prime}$ so that if two points in $\mathcal{B}^{\prime}_{n}$ are in a box of radius $r^{\prime}$, then their images under $p^{\nu}$ are pushed apart by at least $1+\lambda^{\prime}$. By taking $r^{\prime}$ slightly smaller, we may assume it’s a dyadic rational. Now set
@@
This implies $\beta\leq r^{\prime}<2r^{\prime}$ and $\varepsilon_{n+t}+2\beta\leq b$ for all $t\geq 0$. We further have $\varepsilon_{n+t}+2\beta+2\delta<4\delta+\varepsilon_{n}$. for all $t\geq 0$. This follows since $\varepsilon_{n+t}\leq\varepsilon_{n}$ (because the box diameter for boxes in a $\mathcal{B}_{k}$ is $\varepsilon_{k}=2R/2^{k}$ where $R$ is determined by $p$) and $\beta<\delta$. Hence the boxes of side-length $2\beta+\varepsilon_{n+t}$ (with the same centers as the boxes in $\mathcal{B}^{\prime}_{n+t}$) have at least a radius of $\delta$ which is smaller than the boxes in $\mathcal{N}_{1}$ as these boxes have side-length $4\delta+\varepsilon_{n}$. Hence, the distance between $\mathcal{N}_{1}$ and any set of boxes formed from $\beta$-neighborhoods of any refined collection $\mathcal{B}^{\prime}_{n+t}$, and these $\beta$-neighborhoods lie within the region on which $|Dp^{\nu}|>L=1+\lambda^{\prime}/2$.
Given any positive integer $t$, we may consider the refined collection $\mathcal{B}^{\prime}_{n+t}$: subdivide the boxes of $\mathcal{B}^{\prime}_{n}$ by placing a $(2^{t})^{2}$-grid in each box of $\mathcal{B}^{\prime}_{n}$, then repeat the Algorithm 3.2, building edges where a box’s (approximate but error-controlled) image overlaps another box, then refining to yield only boxes that lie in cycles and selecting only (but all of) the expanding components. Whenever we refine, since the $\mathcal{B}_{k}$ sequence is nested, the smaller boxes are in one of the prior boxes so the expansion inequality doesn’t change, and we have expansion by $1+\lambda^{\prime}$ on $\mathcal{B}^{\prime}_{n+t}$.
Then we may consider the set of boxes with the same centers as the centers of the boxes in $\mathcal{B}^{\prime}_{n+t}$, but with side-length increased by $2\beta$, to $\varepsilon_{n+t}+2\beta$; this is $\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)=\cup\{\mathcal{N}(B_{j},\beta):{B_{j}\in\mathcal{B}^{\prime}_{n+t}}\}.$
The choice of $\beta$ yields that for $t\geq 0$, as was stated above, the boxes in $\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$ lie within the neighborhood $\mathcal{N}_{1}$, in fact in a subset with Hausdorff distance $\delta$ to $\mathcal{N}_{1}$. Since on $\mathcal{N}_{1}$, $p^{\nu}$ is expanding by $L$, the same holds on subsets. Thus for any $t\geq 0$, (iii) holds for $\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$ and note $\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)\subset\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta+\delta)\subset\mathcal{N}_{1}$, which helps with (iv).
On the other hand, we have $\mathcal{B}^{\prime}_{k}\subset\mathcal{R}(\alpha_{k})$, where $\alpha_{k}\downarrow 0$ as $k\to\infty$ and is computable. We use shadowing to find how close a real orbit is to our pseudo-orbits. Since $p^{\nu}$ is hyperbolic (because $p$ is), it has the shadowing property on each chain component. We apply Equation (4.29) in [13] to our setting. This provides a minimum (denoted by $\xi$ in [13]) which in our setting is $2r^{\prime}$ because the largest ball that fits inside of an image of a ball of radius $2r^{\prime}$ has a radius bounded by $\lvert Dp^{\nu}\rvert 2r^{\prime}>2r^{\prime}$ since $\lvert Dp^{\nu}\rvert>1$. By Proposition 4.3.4 in [13], we deduce that since $\beta<2r^{\prime}$ (by choice of $\beta$ above), we can compute a $t^{\prime}$ sufficiently large that
@@
Hence $\mathcal{B}^{\prime}_{n+t}$ lies within a $\beta$-neighborhood of $J$, i.e., $\mathcal{B}^{\prime}_{n+t}\subset\mathcal{N}(J,\beta)$. Thus, each box in $\cup_{B_{j}\in\mathcal{B}^{\prime}_{n+t}}\mathcal{N}(B_{j},\beta)$ contains a point of $J$, for all $t\geq t^{\prime}$.
But we also know $J\subset\mathcal{B}^{\prime}_{n+t}\subset\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$. So, we have (i) $J\subset\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$ and (ii) each box in $\cup\{\mathcal{N}(B_{j},\beta):{B_{j}\in\mathcal{B}^{\prime}_{n+t}}\}$ contains a point of $J$ for any $t\geq t^{\prime}$.
@@
It suffices to choose a $t\geq t^{\prime}$ and set $\mathcal{N}_{2}:=\cup\{\mathcal{N}(B_{j},\beta):B_{j}\in\mathcal{B}^{\prime}_{n+t}\}=\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$, so $\beta^{\prime}:=\varepsilon_{n+t}+2\beta.$ By definition of $\beta$ we then have $\beta^{\prime}\in(0,b]$ as desired, (i),(ii),(iii) hold, and $\mathcal{N}(\mathcal{N}_{2},\delta)=\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta+\delta)\subset\mathcal{N}_{1}$ as was explained a few paragraphs above, so (iv) holds.

### v2

@@ Appendix A More Proof Details
For completeness, we include here more details for this proof.
First, given any positive integer $t$, we may consider the refined set $\mathcal{B}^{\prime}_{n+t}$: subdivide the boxes of $\mathcal{B}^{\prime}_{n}$ by placing a $(2^{t})^{2}$-grid in each box of $\mathcal{B}^{\prime}_{n}$, then repeat the Algorithm 3.2, building edges where a box’s (approximate but error-controlled) image overlaps another box, then refining to yield only boxes that lie in cycles and selecting only (but all of) the expanding components. Whenever we refine, since the $\mathcal{B}_{k}$ sequence is nested, the smaller boxes are in one of the prior boxes so the expansion inequality doesn’t change, so we have expansion by $1+\lambda^{\prime}$ on $\mathcal{B}^{\prime}_{n+t}$.
Now, due to this expansion, applying Corollary 2.19 of [3] to $p^{\nu}$ on $\mathcal{B}^{\prime}_{n}$, we compute a radius $r^{\prime}$ so that if two points in $\mathcal{B}^{\prime}_{n}$ are in a box of radius $\leq r^{\prime}$, then the distance between their images under $p^{\nu}$ are pushed apart by at least $1+\lambda^{\prime}$ times the distance between the two original points. By taking $r^{\prime}$ slightly smaller, we may assume it’s a dyadic rational. Now set
@@
This implies $\beta\leq r^{\prime}<2r^{\prime}$ and $\varepsilon_{n+t}+2\beta\leq b$ for all $t\geq 0$, because $\varepsilon_{n+t}\leq\varepsilon_{n}$ (because the box diameter for boxes in a $\mathcal{B}_{k}$ is $\varepsilon_{k}=2R/2^{k}$ where $R$ is determined by $p$). We will set $\beta^{\prime}=\varepsilon_{n+t}+2\beta$ for a $t$ to be calculated near the end of this proof, so we will have our $\mathcal{N}_{2}$ be the union of boxes with centers in some $\mathcal{B}^{\prime}_{n+t}$ but with sidelength $\beta^{\prime}$, hence radii $\varepsilon_{n+t}/2+\beta$. Thus, we will have $\mathcal{N}_{2}\subset\mathcal{N}(\mathcal{B}^{\prime}_{n},\beta)$. Now, the distance from the boundary of $\mathcal{B}^{\prime}_{n+t}$ to the boundary of $\mathcal{N}_{1}$ is at least $2\delta$, because $d_{H}(\mathcal{B}_{n},\mathcal{N}_{1})=2\delta$ and $\mathcal{B}_{n+t}\subset\mathcal{B}_{n}$. Since $\beta<\delta,$ we have the $\beta$-neighborhood of any box in $\mathcal{B}^{\prime}_{n+t}$ is contained in the $\delta$-neighborhood of $\mathcal{B}^{\prime}_{n}$, and hence at least a distance $\delta$ to $\mathcal{N}_{1}$. Thus, we have both $d_{H}(\mathcal{N}_{2},\mathcal{N}_{1})>\delta$ and $\mathcal{N}(\mathcal{N}_{2},\delta)\subset\mathcal{N}_{1}$, which will provide (iv).
Since on $\mathcal{N}_{1}$, $p^{\nu}$ is expanding by $L$, the same holds on subsets of $\mathcal{N}_{1}$. Thus for any $t\geq 0$, (iii) holds for $\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$ hence will hold for our $\mathcal{N}_{2}$.
Now we turn to finding the large enough $t$. We have $\mathcal{B}^{\prime}_{k}\subset\mathcal{R}(\alpha_{k})$, where $\alpha_{k}\downarrow 0$ as $k\to\infty$ and is computable. We use shadowing to find how close a real orbit is to our pseudo-orbits. Since $p^{\nu}$ is hyperbolic (because $p$ is), it has the shadowing property on each chain component. We apply Equation (4.29) in [13] to our setting. This provides a minimum (denoted by $\xi$ in [13]) which in our setting is $2r^{\prime}$ because the largest ball that fits inside of an image of a ball of radius $2r^{\prime}$ has a radius bounded by $\lvert Dp^{\nu}\rvert 2r^{\prime}>2r^{\prime}$ since $\lvert Dp^{\nu}\rvert>1$. By Proposition 4.3.4 in [13], we deduce that since $\beta<2r^{\prime}$ (by choice of $\beta$ above), we can compute a $t^{\prime}$ sufficiently large that
@@
Hence $\mathcal{B}^{\prime}_{n+t}$ lies within a $\beta$-neighborhood of $J$, i.e., $\mathcal{B}^{\prime}_{n+t}\subset\mathcal{N}(J,\beta)$. Thus, if (consistent with earlier notation) $\mathscr{B}^{\prime}_{n+t}$ is our collection of boxes whose union defines $\mathcal{B}^{\prime}_{n+t}$, then each box in $\cup_{B_{j}\in\mathscr{B}^{\prime}_{n+t}}\mathcal{N}(B_{j},\beta)$ contains a point of $J$, for all $t\geq t^{\prime}$.
But we also know $J\subset\mathcal{B}^{\prime}_{n+t}\subset\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$. So, we have (i) $J\subset\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$ and (ii) each box in $\cup\{\mathcal{N}(B_{j},\beta):{B_{j}\in\mathscr{B}^{\prime}_{n+t}}\}$ contains a point of $J$ for any $t\geq t^{\prime}$.
@@
It suffices to choose a $t\geq t^{\prime}$ and set $\mathcal{N}_{2}:=\cup\{\mathcal{N}(B_{j},\beta):B_{j}\in\mathscr{B}^{\prime}_{n+t}\}=\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta)$, so $\beta^{\prime}:=\varepsilon_{n+t}+2\beta.$ By definition of $\beta$ we then have $\beta^{\prime}\in(0,b]$ as desired, (i),(ii),(iii) hold, and $\mathcal{N}(\mathcal{N}_{2},\delta)=\mathcal{N}(\mathcal{B}^{\prime}_{n+t},\beta+\delta)\subset\mathcal{N}_{1}$ as was explained a few paragraphs above, so (iv) holds.